Kyd and North are playing a game. Kyd selects one card from a standard 52-card deck. If Kyd selects a face card (Jack, Queen, or King), North pays him $6. If Kyd selects any other type of card, he pays North $2.


1
Expert's answer
2021-02-24T06:55:05-0500

There are 12 face cards in the standard card-deck of 52 cards.

Therefore, the probability to take face card is equal to 12/52=3/13.

The probability to take another card is equal to 1-3/13=10/13.

Kid's average win is $6*3/13 - $2*10/13 = - $2/13.

This game in average is not profitable for him.